a) Verify that the total differential equation
yz dx +(x^2y-zx)dy+(x^2z-xy)dz=0 is integrable and hence find its integral.
1
Expert's answer
2020-04-03T18:19:50-0400

The given differential equation is ,

"yzdx+(x^2y-zx)dy+(x^2z-xy)dz=0"

We known that ,the necessary and sufficient condition of integrability of the given differential equation "Pdx+Qdy+Rdz=0" is



"P(\frac{\partial Q}{\partial z}-\frac{\partial R}{\partial y})+Q(\frac{\partial R}{\partial x}-\frac{\partial P}{\partial z} )+R(\frac{\partial P}{\partial y}-\frac{\partial Q}{\partial x})=0"

"...........(1)"

Now ,according to question,

"P=yz,Q=(x^2y-zx) \ and \ R=(x^2z-xy)"

Now ,

"\frac{\partial P}{\partial y}=z \ and \ \frac{\partial P}{\partial z}=y.""\frac{\partial Q}{\partial x}=(2xy-z) \ and \ \frac{\partial Q}{\partial z}=-x""\frac{\partial R}{\partial x}=(2xz-y) \ and \ \frac{\partial R }{\partial y}=-x"



putting above all value in equation (1) ,we get ,

"yz(-x+x)+(x^2y-zx)\{(2xz-y)-(y) \}+(x^2z-xy) \{(z)-(2xy-z)\}"

"=0+2(x^3yz-x^2y^2-x^2z^2+xyz+x^2z^2-x^3yz-xyz+x^2y^2)=0"

Hence ,given differential equation is integrable.

To solve the given differential equation , we take

"x=constant. \ \implies \ dx=0" .

"\therefore \ (x^2y-zx) dy+(x^2z-xy)dz=0"

"\implies \ x^2(ydy+zdz)-x(zdy+ydz)=0"

"\implies \ x^2(ydy+zdz)-xd(yz)=0"

On integrating ,we get


"\frac{x^2 y^2} {2}+\frac{x^2z^2}{2}-xyz=\phi \ ........(2)"

Where "\phi" is a constant, it may be regarded as a function of "x."

Differentiating (2) w.r.t. x , y, and z we get,

"(xy^2dx+yx^2dy)+(xz^2dx+zx^2dz)-(yzdx+xzdy+xydz)=d\phi"

On simplifying, we get

"(yx^2-xz)dy+(zx^2-xy)dz+yzdx+(xy^2+xz^2-2yz)dx=d\phi \ ...........(3)"

Since ,"yzdx+(yx^2-xz)dy+(zx^2-xy)dz=0 \ (given \ equation)" ,therefore equation (3) is,

"(xy^2+xz^2-2yz)dx=d\phi \ ........(4)"

Again from equation (2),


"\frac{x^2y^2}{2}+\frac{x^2z^2}{2}-xyz=\phi"

"\implies \ x^2y^2+x^2z^2-2xyz=2\phi"

Now dividing both side by x ,we get


"\implies \ xy^2+xz^2-2yz=\frac{2\phi}{x}............(5)"

From equation (4) and (5) ,we get



"\frac{2\phi }{x} dx=d\phi""\implies \ \frac{2dx}{x}=\frac{d\phi }{\phi}"

On integrating ,we get

"logx^2+logk=log\phi"

Where k is a constant.

"\implies \ log(x^2k)=log\phi"

"\implies \ x^2k=\phi"

Now ,putting the value of "\phi" in equation (2) ,we get


"\frac{x^2y^2}{2}+\frac{x^2z^2}{2}-xyz=x^2k"

"\implies \ xy^2+xz^2-2yz=2xk"